OpenStudy (anonymous):

Anyone in Algebra 2?!?!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

*doing moonwalk*

OpenStudy (anonymous):

i really need help!

OpenStudy (anonymous):

okay which one I need the Unit, lesson

OpenStudy (anonymous):

i private messaged u @Jelia97

OpenStudy (anonymous):

private msgs for me isnt working today

OpenStudy (anonymous):

idk y, but give me the unit and lesson

OpenStudy (anonymous):

ohh ok. unit 7, lesson 6 and 8

OpenStudy (anonymous):

also the semester unit exam

OpenStudy (anonymous):

oh i have unit 7 lesson 6 done i still need the test too

OpenStudy (anonymous):

ok can u give me that one please?!?

OpenStudy (anonymous):

1. C log8a-log82 2. D 2.089 3. A 3.57 4. A 0.82 5. D 41.67 6. B In x^2/c^5 7. A 21,405 mowers 8. log3(40)-log3(10) log3(40/10) log3(4) answer 9. log (1/16)=(4*x-3)* log 64 - log 16 = (4*x - 3)*log 64 4*x - 3 = -(log 16/log 64) = - log 2^4 / log 2^6

OpenStudy (anonymous):

10. 3 log(2x) = 4 log(2x)=4/3 10^(4/3)=2x 2x≈21.5443 x≈21.5443/2 x≈10.7722

OpenStudy (anonymous):

thank u sooo much!!!!! you,re a lifesaver!!!

OpenStudy (anonymous):

*you're haha

OpenStudy (anonymous):

11. e^8=5x+7 5x=e^8-7 x=(e^8-7)/5 x≈594.792

OpenStudy (anonymous):

so these are the right ones right?

OpenStudy (anonymous):

#12 i got partial credit on so idk if you want that one or not?

OpenStudy (anonymous):

yeah everything else is perfect!

OpenStudy (anonymous):

yeah can i get that one too(:

OpenStudy (anonymous):

12. 5e^(2x)=19 e^(2x)=19/5 2x=ln(19/5) x=0.668

OpenStudy (anonymous):

THANK U SO MUCH!!

OpenStudy (anonymous):

ur welcome!

OpenStudy (anonymous):

if u get unit test or semester test answers can u get them too me? that would be amazing! and if i get them before u do i will get them to u!